# Anomalous Criticality of Absorbing State Transition toward Jamming

### v3

@@ IV Results
In the Griffiths phase, the quenched disorder generates strong dynamic heterogeneity in the relaxation process. This can be visualized by the spatial distribution of normalized activity duration $f_{ad}$, i.e., the fraction of time that a particle stays in an active state relative to the active duration of the whole system, starting from a perturbed absorbing configuration. In Fig. 3(d), we show these distributions for systems with $\varepsilon=0.2\sigma$. One can observe non-uniform network structures of activity, resembling the dynamic heterogeneity of supercooled liquids or the force chain of the jammed state [peters2005characterization, radjai1996force, majmudar2007jamming]. In Regime II, one can also find relatively weak dynamic heterogeneity (Fig. S8), which would be smoothed out by further increasing $\varepsilon$ to Regime I.
In the Griffiths phase, the finite-size effect of $f_{b}(t)$ and $P_{s}(t)$ near criticality is also reversed, i.e., at the same density near the transition, small systems stay longer in the active state than large systems. This is best shown in Fig. 3(b) and Fig. 4(a), as well as in Fig. S2. This abnormality is nevertheless consistent with the fact that small systems are more likely to be jammed near the jamming point [goodrich2012finite, xu2005random]. This suggests that close-packing significantly influences the criticality of absorbing phase transition in this regime. Further support comes from the structure factor $S(q)$ for systems with $\varepsilon\rightarrow 0$, which does not follow the critical $q^{0.45}$ scaling of the Manna model, as expected by the field theory (see Fig. S9).
To further prove that the above phenomena are results of the Griffiths effect, we manually add quenched disorder in systems located at Regime I, by making $\varepsilon$ spatially dependent. More explicitly, the system is divided into cells of size $l$. Each cell $(i,j)$ has its own $\varepsilon_{i,j}$ set as
@@
Figure 4: (a, c) Finite size effects of survival probability of dynamic trajectory $P_{s}(t)$. (d, f) Distributions of normalized activity duration. (a, b) are for the original BRO models near the transition point at $\varepsilon=0.01\sigma$ and $\varphi=0.8380$. (c, d) are for the heterogeneous contact model at $P_{e}=0.788$ and $k=1$ based on the same configuration as in (b).
@@
Finally, we study the critical behaviors of the BRO model in the close packing limit ($\varepsilon\rightarrow 0$). In this limit, the particle configuration can be regarded as static during the relaxation process. From the perspective of field theory, this scenario corresponds to $\theta=0$ in Eq.(4), where the evolution of density field $\psi$ is completely forbidden. Thus, $\psi$ acts as pure quenched disorder in the dynamics of $\rho$ field, essentially changing the CDP universality to the directed percolation (DP) with quenched disorder [hooyberghs2004absorbing, moreira1996critical, vojta2005critical]. The classical contact model is the simplest realization of DP [1974Contact, jensen1992critical]. To test this idea, we simulate a contact model on the networks composed of nearly jammed particles generated by the BRO model at $\varepsilon=0.01\sigma$, where each particle represents a node. We find that the topology of the contact network alone does not affect the critical behaviors of systems. However, if the infection rate of a node is proportional to the activity durations of the corresponding particle (see the Method section for details), the activity distribution of this contact model exhibits dynamic heterogeneity resembling that of the high density BRO model, as shown in Fig. 4(b, d). Correspondingly, the finite size scaling of $P_{s}(t)$ in Fig. 4(c) also exhibits similar anomalous behaviors as the BRO model in the close packing limit in Fig. 4(a). This similarity proves that the critical behaviors of the BRO model in the nearly jammed configuration correspond to DP on hierarchical heterogeneous network. Moreover, the Griffiths effect becomes stronger as the heterogeneity increases in the DP model, i.e., for $k=2,4$. (see Fig. S10). It should be noted that our previous modified BRO model with cell-size quenched-disorder fails to account for the Griffiths effect at the close packing limit, because the correlation at this point spans the system size, leaving the topology of heterogeneous network to play the dominant role.
V Conclusion
@@
In summary, we investigate the critical behavior of the BRO model in dense monodisperse and binary systems in $d=2$ to 4. We find that anomalous critical phenomena emerge as the critical density increases by decreasing $\varepsilon$. For 3D monodispersed systems at intermediate critical density, the crystallization interrupts the dynamic phase transition. For binary systems at intermediate critical density, the measured critical exponents deviate from the Manna universality class due to the caging effects, signifying the existence of new universality of dynamic phase transition, i.e., an absorbing to active glass transition. For both monodispersed and binary systems at higher density, the quenched disorder of contact networks triggers the pronounced Griffiths effect, smearing out the dynamic phase transition. In the $\varepsilon\rightarrow 0$ limit, the BRO model becomes equivalent to DP on heterogeneous network. Our results demonstrate that although the BRO model can generate RCP configurations, the criticality of the dynamic transition is fundamentally reshaped by the geometric rigidity of the RCP state. These findings discourage previous attempts to directly associate the jamming transition with the Manna universality class. We also propose a field theory with fractional time dynamics which unifies both the absorbing to active-glass transition and heterogeneous DP, revealing a deep connection between dynamic criticality and spatial annealed/quenched disorder. Our work could potentially be related to ultra-stable glass and the Gardner transition [wang2025hyperuniform, 2015Growing, charbonneau2015numerical, wang2024gardner, anand2025emergent] and opens the door to studying the dynamic criticality of more complicated artificial neural networks [spigler2018jamming], like the learning process of spherical negative perceptron model [zhang2024absorbing].
@@ VI Method
For typical second-order absorbing phase transitions, $f_{a}(t)$ should satisfy the finite-size scaling relationship [lubeck2003universal, zheng1996generalized, lei2021barrier]
@@
where $\mathcal{P}(x,0)\sim 1$ for $x\ll 1$ and $\mathcal{P}(x,0)\rightarrow 0$ for $x\gg 1$. From Eq. (10), we can also obtain the finite size scaling relationship for the characteristic time of the active spreading process $t^{*}$ satisfying $\mathcal{P}(t^{*}L^{-z^{*}},\tau L^{1/\nu_{\perp}})=0.5$, i.e.,
@@
In the classical contact model, two important controlling parameters are the basic infection rate $P_{e}$ and recovery rate $P_{r}$. The probability that the healthy site $i$ becomes infected by an adjacent infected site in the next time step is
@@
Here $f_{ad}(i)$ is in the range of $(0,1)$. As $k$ increases, the heterogeneity becomes more prominent. We study different basic infection rate cases under kk=1, 2 and 4. We find that the Griffith effect becomes stronger as $k$ increases (see Fig. S10).
@@ Acknowledgments:
The authors are grateful to Ran Ni, Ning Xu, Hao Hu and Hua Tong for helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12347102, No. 12275127), the Innovation Program for Quantum Science and Technology (No. 2024ZD0300101), the Natural Science Foundation of Jiangsu Province (No. BK20250058, No. BK20233001), Program for Innovative Talents and Entrepreneurs in Jiangsu, the Fundamental Research Funds for the Central Universities (0204-14380249, KG202501). The simulations are performed on the High-Performance Computing Center of Collaborative Innovation Center of Advanced Microstructures, the High-Performance Computing Center (HPCC) of Nanjing University.

### v4

@@ IV Results
In the Griffiths phase, the quenched disorder generates strong dynamic heterogeneity in the relaxation process, characterized by quasi-static network composed of frequently overlapping particles. This can be visualized by the spatial distribution of normalized activity duration $f_{ad}$, i.e., the fraction of time that a particle stays in an active state relative to twice of the average active duration of the whole system (See the SI for details). In Fig. 5(c), we show these distributions for systems with $\varepsilon=0.1\sigma$. One can observe non-uniform network structures of activity, resembling the dynamic heterogeneity of supercooled liquids or the force chain of the jammed state [50, 54, 56, 40]. In Regime II, one can also find relatively weak dynamic heterogeneity, which would be smoothed out by further increasing $\varepsilon$ to Regime I (Fig. S13).
In the Griffiths phase, the finite-size effect of $f_{b}(t)$ and $P_{s}(t)$ near criticality is also reversed, i.e., at the same density near the transition, small systems stay longer in the active state than large systems. This is best shown in Fig. 5(b) and Fig. 6(a). This abnormality is nevertheless consistent with the fact that small systems are more likely to be jammed near the jamming point [21, 86]. This suggests that close-packing significantly influences the criticality of absorbing phase transition in this regime.
To further prove that the above phenomena are results of the Griffiths effects, we manually add quenched disorder in systems located at Regime I, by making $\varepsilon$ spatially dependent. More explicitly, the system is divided into cells of size $l$. Each cell $(i,j)$ has its own $\varepsilon_{i,j}$ set as
@@
Figure 6: (a, c) Finite size effects of survival probability of dynamic trajectory $P_{s}(t)$. (b, d) Distributions of normalized activity duration. (a, b) are for the original BRO models near the transition point at $\varepsilon=0.01\sigma$ and $\varphi=0.8380$. (c, d) are for the heterogeneous contact model at $P_{e}=0.780$ and $k=1$ based on the same configuration as in (b).
@@
Finally, we study the critical behaviors of the BRO model in the close packed limit ($\varepsilon\rightarrow 0$). In this limit, the particle configuration can be regarded as static during the relaxation process (corresponding to $\theta\rightarrow 0$ in Eq. (4)). Although the spatial fluctuation of $\psi$ field is small, it acts as strong quenched disorder influencing the dynamics of $\rho(t)$, as the coupling coefficient $\omega$ diverges in the close-packed limit.
More explicitly, the fluctuation of density field around mean value $\psi_{0}$ can be decomposed into plastic contribution $\delta\psi_{p}(t)$ and elastic contribution $\delta\psi_{e}(t)$ [73, 34], i.e.,
$\delta\psi(t)=\delta\psi_{p}(t)+\delta\psi_{e}(t)$ (8)
where $\delta\psi_{p}(t)$ is caused by the rearrangement of the topological networks of particles due to cage-breaking relaxation, while $\psi_{e}(t)$ is attributed to vibrational displacements of particles in the cage, whose magnitude can be reflected by Debye-Waller factor [34]. For disordered systems near jamming point, $\psi_{p}(t)$ freezes and hardly evolves, becoming a spatial quenched disorder $\delta\psi_{qc}$ whose influence is persistent and much stronger than the fluctuating $\delta\psi_{e}(t)$. By neglecting $\delta\psi_{e}(t)$ field, Eqs. (3-4) reduce to the field theory of directed percolation (DP) with quenched disorder, or heterogeneous DP  [31, 46, 76].
$\frac{\partial\rho(t)}{\partial t}$ $=$ $D\nabla^{2}\rho+a^{\prime}\rho-b\rho^{2}+\omega\rho\delta\psi_{\mathrm{qc}}+\sigma_{n}\sqrt{\rho}~\xi(t)$~~~ (9)
@@
where $a^{\prime}=a+\omega\psi_{0}$. The classical contact model on a hierarchical heterogeneous network is the simplest realization of heterogeneous DP [24, 32, 77]. To test the above idea, we simulate a contact model on the networks composed of nearly jammed particles generated by the BRO model at $\varepsilon=0.01\sigma$, where each particle represents a node. We find that the topology of the contact network alone does not affect the critical behaviors of systems. However, if the infection rate of a node is proportional to the activity durations of the corresponding particle, i.e., $k=1$ (see the Method Section for details), the activity duration distribution of this contact model exhibits dynamic heterogeneity resembling that of the high density BRO model(Fig. 6(b, d)). Correspondingly, the finite size scaling of $P_{s}(t)$ in Fig. 6(c) also exhibits similar anomalous behaviors as the BRO model in the close packing limit ( Fig. 6(a)). Moreover, the Griffiths effects become stronger as the heterogeneity increases in the DP model, i.e., for $k=2,4$. (see Fig. S14). These similarities prove that the BRO model in the nearly jammed configuration corresponds to heterogeneous DP model.
$\varepsilon$ $\alpha$ $\nu_{\perp}$ $\nu_{\parallel}$ $\beta$ $^{*} zz$
[…]
Manna 0.419 0.799 1.225 0.639 1.533 1.533
Table 1: Critical exponents measured in 2D crystal systems under conserved BRO dynamics. Reference values of the Manna universality class are from Ref. [39, 25].
Dynamic criticality in crystal phase
In a previous study, the absorbing transition of conserved BRO model in crystal phase has also been suggested to reside in the Manna universality [19]. We analyse this issue within the above theoretical framework. For crystal structures without defects, the quenched disorder disappears, i.e., $\delta\psi_{p}(t)=0$. Thus Griffiths effect is absent and the fluctuation of $\psi(t)$ is solely determined by $\delta\psi_{e}(t)$ which is connected to local strain (displacement) field ${\mathbf{u}}(t)$, i.e., $\psi_{e}(t)=\psi_{0}\nabla{\mathbf{u}}(t)$ [10]. The particle velocity field is $\bm{v}(t)={\partial{\mathbf{u}}(t)}/{\partial t}$. This gives the field equation for critical dynamics of crystal phase
$\frac{\partial\rho(t)}{\partial t}=D\nabla^{2}\rho+a\rho-b\rho^{2}+\omega\rho\delta\psi_{e}(t)+\sigma_{n}\sqrt{\rho}~\xi(t)$~~~~~~ (10)
$\frac{\partial\delta\psi_{e}(t)}{\partial t}=\psi_{0}\nabla{\bm{v}}(t),$ (11)
In isotropic systems, $\bm{v}(t)=\Gamma\nabla p(t)$ where $p(t)$ is the effective pressure. Near the critical point, $p(t)$ should be proportional to $\rho(t)$ as a first order approximation, which gives $\bm{v}(t)\sim\nabla\rho(t)$ [62, 53]. Then Eq. (10-11) returns to the field equation for the Manna universality class. However, the intrinsic anisotropy of crystal structures does not guarantee this reduction. In Table I and Fig. S16, we give the measured critical exponents for crystal phase. We find notable deviations from the Manna universality class for dynamic critical exponents $\alpha$, $\nu_{\parallel}$ and $z$. The difference between our results and Ref. [19] might be that Ref. [19] did not use defectless crystal structures as initial configurations.
Figure 7: Structure factor for 2D conserved binary BRO systems. For (b), $\varphi_{c}=0.693$.
Long wavelength density fluctuations
Finally, we discuss the density fluctuation of the system. It is known that at the critical point of RO/BRO model at large $\varepsilon$, the density fluctuation is suppressed, giving rise to the critical hyperuniform state with structure factor $S(q)\sim q^{0.45}$ (2D) and $S(q)\sim q^{0.25}$ (3D) [71, 26]. This critical hyperuniform state is a dynamic state accompanied by the intermittent diffusion of active particles. It was argued that this hyperuniform scaling remains unchanged when the system approaches the jammed point $\varepsilon\rightarrow 0$ [84, 85]. Nevertheless, the exponent 0.25 for 3D system is too small to be accurately determined near the jamming point (Fig. S18). For 2D conserved binary BRO model, we find similar hyperuniformity scaling $S(q)\sim q^{0.45}$ when $\varepsilon=0.6\sigma$ (Fig. 7(b)), consistent with Ref. [84]. However, by further decreasing $\varepsilon$ to $0.1\sigma$, an obvious deviation of $S(q)$ from $q^{0.45}$ can be found (Fig. 7(a)). Similar behaviors can also be found in non-conserved BRO model in Fig. S17. This indicates that $S(q)$ at jamming point is controlled by the contact network of jamming configuration, rather than the critical dynamics of the Manna universality class.
V Discussion and Conclusion
In the above analysis, we only discuss the conserved BRO model. Actually, the non-conserved BRO model exhibits qualitatively the same results(Fig. S1, S2, S6, S8, S10, S11, S15, S17, Table SI). Quantitatively, the non-conserved BRO model introduces additional randomness compared with the conserved BRO model. Thus the caging-effect is relatively weaker and the active glass phase area is smaller than that of the conserved BRO model (Fig. S11).
@@ V Discussion and Conclusion
In summary, we investigate the critical behavior of two types of BRO model in dense monodisperse and binary systems in $d=2$ to 4. We find that anomalous critical phenomena emerge as the critical density increases by decreasing displacement size $\varepsilon$. For 3D monodisperse systems at intermediate critical density, the crystallization interrupts the dynamic phase transition. For binary systems at intermediate critical density, the measured critical exponents deviate from the Manna universality class due to the caging effects, signifying the existence of a new universality class of dynamic phase transition, i.e., the absorbing to active glass transition. For disordered systems at higher density, the quenched disorder of contact networks triggers the pronounced Griffiths effects, smearing out the dynamic critical point. In the $\varepsilon\rightarrow 0$ limit, the BRO model becomes equivalent to DP on a heterogeneous network. Our results demonstrate that although the BRO model can generate RCP configurations, the criticality of the dynamic transition is fundamentally reshaped by the geometric rigidity of the RCP state. These findings discourage previous attempts to directly associate the jamming transition in disordered systems with the Manna universality class under $d\leq 4$. For close-packed crystal phase, the Griffiths effect is absent, but we find that the dynamic criticality still deviates from the Manna universality class due to the anisotropy of the crystal structures. We propose a field theory with fractional time dynamics which unifies the above complicated dynamic behaviors, revealing a deep connection between dynamic criticality, spatial annealed/quenched disorder and jamming. Our work could potentially be related to ultra-stable glass and the Gardner transition [80, 8, 11, 79, 1], and enriches the studies of the dynamic criticality of more complicated artificial neural networks [68, 1], like the learning process of spherical negative perceptron model [87].
@@ VI Method
For typical second-order absorbing phase transitions, $f_{a}(t)$ should satisfy the finite-size scaling relationship [25, 38, 88, 35]
@@
where $\mathcal{P}(x,0)\sim 1$ for $x\ll 1$ and $\mathcal{P}(x,0)\rightarrow 0$ for $x\gg 1$. From Eq. (14), we can also obtain the finite size scaling relationship for the characteristic time of the activity decay $t^{*}$ satisfying $\mathcal{P}(t^{*}L^{-z^{*}},\tau L^{1/\nu_{\perp}})=0.5$, i.e.,
@@
In the classical contact model, two important controlling parameters are the basic infection rate $P_{e}$ and recovery rate $P_{r}$ [24, 32, 25]. The probability that the healthy site $i$ becomes infected by an adjacent infected site in the next time step is
@@
Here $f_{ad}(i)$ is in the range of $(0,1)$. As $k$ increases, the heterogeneity becomes more prominent. We study different basic infection rate cases under kk=1, 2 and 4. We find that the Griffiths effects become stronger as $k$ increases (see Fig. S14).
@@ Acknowledgments:
The authors are grateful to Ran Ni, Ning Xu, Hao Hu and Hua Tong for helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12347102, 12275127), the Natural Science Foundation of Jiangsu Province (No. BK20233001, No. BK20250058), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM502), Program for Innovative Talents and Entrepreneur in Jiangsu, the Fundamental Research Funds for the Central Universities (0204-14380249, KG202501), the Innovation Program for Quantum Science and Technology (No. 2024ZD0300101). The simulations are performed on the High-Performance Computing Center of Collaborative Innovation Center of Advanced Microstructures, the High-Performance Computing Center (HPCC) of Nanjing University.